OpenStudy (anonymous):

ln(x^2+1)=8 solve the logarithmic function algebraically. round the result to three decimal places.

myininaya (myininaya):

\[e^{\ln(x^2+1)}=e^8\]

myininaya (myininaya):

\[x^2+1=e^8\]